OpenStudy (anonymous):

solve by elimination method x + 4y = 9 -x + 4y = -1 what is the solution of system

OpenStudy (anonymous):

help someone please solve by elimination method x + 4y = 9 -x + 4y = -1 what is the solution of system

OpenStudy (anonymous):

Call x + 4y = 9 equation 1, and -x + 4y = -1 equation 2. You then subtract equation 2 from equation 1 so you get, x - (-x) + 4y - 4y = 9 - (-1) which makes 2x = 10 so :. x = 5. Substitute 5 into equation 1, so it becomes, 5 + 4y = 9 It then becomes 4y = 9 - 5 = 4 :. 4y = 4 and so y = 1, x = 5

OpenStudy (anonymous):

so the solutio is (1, 5) correct